# SceneGen: Single-Image 3D Scene Generation in One Feedforward Pass

### v1

@@ 3.4 Extension to Multi-view Inputs
Despite only trained exclusively on single-image samples, our model surprisingly demonstrates inherent multi-view compatibility through its flexible feature extraction and conditioning strategy. Specifically, for a scene with $K$ input views ($\{\boldsymbol{I}_{\mathrm{Scene}}^{k}\}_{k=1}^{K}$), each view’s visual features ($\mathcal{F}_{V}^{k}$) are independently extracted via the visual encoder ($\Phi_{V}$), while the geometric features are obtained from the unified scene representations encoded by aggregating information across all views with the geometric encoder ($\Phi_{G}$), denoted as:
$\mathcal{F}_{\mathrm{geo}}^{k}=\Phi_{G}(\{\boldsymbol{I}_{\text{Scene}}^{j}\}^{K}_{j=1})[k]$
The final asset positions are determined by averaging predictions across all views. Experiments show that this multi-view inference scheme improves asset generation quality through better geometric understanding, although the model has never been explicitly fine-tuned on such inputs.
@@ 4 Experiments
This section starts with the experimental settings in Sec. 4.1, followed by comprehensive quantitative and qualitative evaluations in Sec. 4.2 and Sec. 4.3, respectively. Finally, we conduct ablation studies in Sec. 4.4 to validate the effectiveness of our proposed model and strategies.
@@ 4.1 Experimental Settings
For visual quality, we focus on the scene texture rendering. Concretely, after alignment with the ground truth point cloud, we render the predicted scenes with Blender from the original input camera viewpoint. We consider two types of ground truth: (i) instance-masked scene images extracted using corresponding object masks, and (ii) images rendered from ground truth assets at the same viewpoint (excluding ambient lighting). We compare our rendered results with both types of ground truth using PSNR, SSIM, LPIPS [70], FID [21], CLIP [44] similarity, and DINOv2 [42] similarity to assess the texture quality of generated assets. Regarding efficiency, we report the inference time cost for synthesizing a single 3D asset on a single A100 GPU. More details will be included in Sec. C.2 of the Appendix.
Baselines. We compare SceneGen with representative 3D scene generation methods, including PartCrafter [34], DepR [72], Gen3DSR [10], and MIDI [23], using their pre-trained models. Specifically, we adopt object masks to specify generation targets for all baselines except for PartCrafter, which does not support mask-based control. Instead, we directly provide PartCrater with extracted objects and the number of assets as input. Moreover, as PartCrafter and DepR do not offer code for texture rendering, our evaluation of these methods focuses on geometric quality, while visual quality is compared with Gen3DSR and MIDI (relying on MV-Adapter [24] for texture synthesis).
@@ 4.2 Quantitative Results
As presented in Table 1, we draw the following key observations: (i) geometric quality: SceneGen consistently outperforms existing methods across all scene-level and asset-level metrics. This stems from its joint integration of local asset features and global scene context during generation. The interactions among multiple assets facilitate the model to produce physically plausible geometric structures, while the position head further improves the structural realism by explicitly predicting spatial arrangements. (ii) visual quality: SceneGen can render high-quality textures for generated 3D assets without relying on any external texture generation models. Moreover, whether using masked scene images or ground-truth renderings as references, our method consistently achieves the best performance across all metrics. This indicates that our synthesized assets are spatially closer to the ground truth while maintaining superior texture fidelity. and (iii) efficiency: While PartCrafter demonstrates a clear advantage in inference speed, it suffers from limited generation quality and controllability. In contrast, SceneGen achieves both superior quality and a strong balance between quality and efficiency, synthesizing a 3D scene containing four assets with geometry and textures within 2 minutes on a single A100 GPU.
In addition, while the baseline methods, e.g., PartCrafter, DepR, and MIDI have been trained on 3D-FRONT [13], which may overlap with our test data, SceneGen still consistently outperforms them across all metrics, further demonstrating its effectiveness and superiority.
@@ 4.3 Qualitative Results
Comparisons with baselines. As depicted in Figure 4, we qualitatively compare SceneGen with existing baselines on both the 3D FUTURE test set and in-the-wild ScnaNet++ [63], where they still struggle with 3D scene generation: PartCrafter lacks controllability over the generated targets and often mistakenly merges distinct assets, while both PartCrafter and DepR are limited to geometry generation and cannot render textures. More critically, all these methods exhibit difficulties in accurately understanding the spatial relationships among assets. In contrast, our proposed SceneGen precisely predicts the spatial relationships among assets and generates multiple 3D assets with accurate geometry and high-quality textures, without relying on any additional tools. These results clearly demonstrate the effectiveness and superiority of our approach.
Extension to multi-image inputs. Benefiting from our architecture design, SceneGen can seamlessly handle multi-image inputs after being trained solely on single-image samples. Given the absence of suitable datasets for quantitative evaluation, we qualitatively assess the impact of multi-image inputs by randomly sampling several scenes from ScanNet++ [63] and employing SAM2 [46] to obtain segmentation masks of corresponding objects. As depicted in Figure 5, compared to single-image inputs, incorporating multi-view images leads to 3D assets with more complete geometry and finer texture details. This illustrates that SceneGen can adaptively integrate complementary information from multiple views to produce higher-quality 3D scenes, further validating its practicality and scalability. More qualitative results will be included in Sec C.1 of the Appendix.
@@ 4.4 Ablation Studies
To validate the effectiveness of our modules, we conduct comprehensive evaluations on several variants of SceneGen, assessing both geometric and visual quality of synthesized scenes. Concretely, we gradually remove global geometric features ($\mathcal{F}_{\mathrm{global}}^{\mathrm{geo}}$), global visual features ($\mathcal{F}_{\mathrm{global}}^{V}$), mask visual features ($\mathcal{F}_{i}^{\mathrm{mask}}$), and substitute the scene-level self-attention block ($\mathcal{A}_{\mathrm{SS}}$) with a simple asset-level self-attention block ($\mathcal{A}_{\mathrm{AS}}$) to evaluate their impact on SceneGen. As depicted in Table 2, we have the following observations: (i) Removing any aforementioned components degrades the overall performance, confirming their importance in SceneGen; (ii) The geometric features primarily affect the structure of synthesized scenes, while the visual features further impact the visual quality; and (iii) The absence of scene-level self-attention blocks eliminates inter-asset interactions during generation, leading to notable performance declines across all metrics. These results strongly demonstrate the necessity and effectiveness of our proposed feature extraction and aggregation modules for SceneGen.
@@ 5 Conclusion
In this paper, we propose SceneGen, which takes a single scene image and target asset masks as input, simultaneously synthesizing multiple 3D assets with structure and texture, as well as relative spatial positions in one feedforward pass. Specifically, we incorporate dedicated visual and geometric encoders to extract both asset-level and scene-level features, which are effectively fused with our introduced feature aggregation module. Notably, through our meticulous design, SceneGen can even directly generalize to multi-image inputs and achieve even better generation quality. Quantitative and qualitative evaluations demonstrate that SceneGen can generate physically plausible and mutually consistent 3D assets, significantly outperforming previous methods in terms of generation quality and efficiency.
@@ Appendix A Preliminaries on 3D Foundation Models
TRELLIS. For a 3D asset ($\mathcal{O}$), TRELLIS encodes its geometry and appearance into a unified representation ($\boldsymbol{z}$), denoted as: $\boldsymbol{z}=\{(\boldsymbol{z}_{i},\boldsymbol{p}_{i})\}_{i=1}^{L}$. Here, $\boldsymbol{p}_{i}\in\{0,1,\ldots,D-1\}^{3}$ denotes the positional index of an active voxel intersecting the surface of $\mathcal{O}$, and $\boldsymbol{z}_{i}\in\mathbb{R}^{C}$ represents the corresponding local latent feature, with $D$ and $L$ representing the 3D grid resolution and the total number of active voxels, respectively.
@@
VGGT. Trained on large-scale 3D annotated data, VGGT can extract 3D scene features through a purely feedforward network without explicit 3D inductive biases. For single or multi-view RGB inputs ($\{\boldsymbol{I}_{i}\}_{i=1}^{s}$), its aggregator derives scene geometric features ($\{\mathcal{F}_{i}^{\mathrm{geo}}\}_{i=1}^{s}$) via:
@@
Here, $\mathcal{F}_{G}^{\mathrm{geo}}$ and $\mathcal{F}_{I}^{\mathrm{geo}}$ represent features extracted by global self-attention and self-attention layers, respectively. These features are efficiently decoded by lightweight DPT layers [45] into depth maps, point maps, and tracks, validating their rich scene geometric representation capacity.
@@ Appendix B More Details about Training Data
We train our SceneGen model on the 3D-FUTURE [14] dataset, whose rich textures and diverse scene lighting conditions effectively simulate real-world environments and thus enhance the model’s generalization ability. In addition, we further scale up training data through data augmentation to ensure the model can robustly learn the relative spatial relationships among multiple assets. Concretely, for a scene with $N$ objects, during training, we iteratively select each asset as the query asset and randomly shuffle the remaining assets. Considering GPU memory constraints, we set the maximum number of assets per scene to $N^{\prime}=7$ on a single A100 GPU. For samples containing more than $N^{\prime}$ assets, we randomly select $N^{\prime}$ assets from them for training. Furthermore, following TRELLIS [59], we also leverage its aesthetic score filtering criterion to filter out assets with aesthetic scores below 4.5, ensuring the high quality of training data. The distribution of asset counts across training scenes is illustrated in Figure 6.
@@ Appendix C More Implementation Details
In this section, we provide a comprehensive explanation of the implementation details discussed in the paper. Specifically, Sec. C.1 describes the specific strategies applied to extend SceneGen to multi-image inputs; and Sec. C.2 elaborates on the details of our evaluation protocols.
@@ C.1 Extension to Multi-image Inputs
While our model is designed for 3D scene generation from a single scene image and is trained only on datasets containing single-view images, SceneGen can be easily adapted to multi-view inputs by simply modifying the sampling process, without requiring additional training or fine-tuning. Concretely, during inference, our model can take images of the same scene from different viewpoints, along with their corresponding objects and instance masks, as input. The geometric encoder ($\Phi_{G}$) within our SceneGen, specifically, an off-the-shelf VGGT [53] aggregator, integrates geometric information from different viewpoints to obtain better geometric representations for each perspective. This enables SceneGen to synthesize better geometric structures during the generation process. Finally, we predict the relative positions among different assets from each viewpoint and use the mean of these predictions across all views as the final spatial position output. It is important to note that, to ensure correctness throughout the inference process, the input order of assets and their segmentation masks must remain consistent across different viewpoints.
@@
Table 3: Geometric Metric Comparisons with Different Point Cloud Alignment Methods.
@@ C.2 Evaluation Protocols
Geometric metrics. Following previous work [23], we conduct geometry evaluation in normalized 3D space (also referred to as canonical space, i.e., $x,y,z\in[-1,1]$), where the ground truth and the synthesized query asset are first rigidly aligned using point cloud registration algorithms. Unlike MIDI [23], which adopts the traditional Iterative Closest Point (ICP [2]) method that produces suboptimal alignment results, we employ FilterReg [17], a fast, accurate, and robust point cloud alignment approach. As presented in Table 3, both MIDI and SceneGen achieve better overall performance when aligned via FilterReg, demonstrating the reliability of this alignment method compared to traditional ICP. Moreover, under both alignment strategies, SceneGen consistently outperforms MIDI, indicating that explicitly predicting the spatial positions among assets enables SceneGen to more accurately model the relationships among distinct 3D assets within the scene.
@@ Appendix D More Visualizations
In this section, we present additional visualization results on the 3D-FUTURE test set to qualitatively compare our SceneGen with representative baselines. As depicted in Figure 8, we have the following observations: (i) PartCrafter frequently suffers from missing or mixed-up assets due to its inability to control generation via object masks, despite already taking segmented objects and asset counts as input; (ii) Both PartCrafter and DepR can only generate scene geometry without rendering texture details; and (iii) All baseline methods (including Gen3DSR and MIDI) share the common limitation of incorrect spatial relationships among synthesized assets. In contrast, our SceneGen fully integrates visual and geometric features within the scene to enable mutual influence among multiple assets during generation, producing 3D scenes with physically 3D scenes and high-quality texture details.
@@ E.1 Limitations
While our SceneGen demonstrates superior performance in 3D scene generation, it is not without its limitations. Specifically, although SceneGen demonstrates better texture generation and generalization capabilities compared to previous methods that rely on canonical representations, the narrow training data distribution limits its ability to generalize to non-indoor scenes, restricting its generalization to a broader range of environments. Moreover, while SceneGen can generate multiple 3D assets and relative spatial positions in a single feedforward pass, without relying on complex post-processing, it does not always handle contact relationships among objects, occasionally leading to asset overlaps or geometric inconsistencies. This is mainly because our single-stage framework does not explicitly enforce strict spatial or physical constraints among objects.

### v2

@@ 3.4 Extension to Multi-view Inputs
Despite being trained exclusively on single-image samples, our model exhibits inherent multi-view compatibility, enabled by its flexible feature extraction and conditioning strategy. Given a scene with $K$ input views ($\{\boldsymbol{I}_{\mathrm{scene}}^{k}\}_{k=1}^{K}$), the visual features ($\mathcal{F}_{V}^{k}$) for each view are extracted independently via the visual encoder ($\Phi_{V}$), while the geometric features are derived from a unified scene representations encoded by aggregating information across all views using the geometric encoder ($\Phi_{G}$), denoted as:
$\mathcal{F}_{\mathrm{geo}}^{k}=\Phi_{G}(\{\boldsymbol{I}_{\mathrm{scene}}^{j}\}^{K}_{j=1})[k]$
The final asset positions are determined by averaging the predictions across all views. Experimental results (detailed in Sec. 4.3) indicate that this multi-view inference scheme improves generation quality by leveraging better geometric understanding, despite the model having never been explicitly fine-tuned on such multi-view inputs.
@@ 4 Experiments
This section starts with the experimental settings in Sec. 4.1, followed by comprehensive quantitative and qualitative evaluations in Sec. 4.2 and Sec. 4.3, respectively. Finally, we conduct ablation studies in Sec. 4.4.
@@ 4.1 Experimental Settings
For visual quality, we focus on the scene texture rendering. Specifically, after alignment with the ground truth point cloud, we render the predicted scenes with Blender from the original input camera viewpoint. We consider two types of ground truth: (i) instance-masked scene images extracted using corresponding object masks, and (ii) images rendered from ground truth assets at the same viewpoint (excluding ambient lighting). We compare our rendered results with both types of ground truth using PSNR, SSIM, LPIPS [70], FID [21], CLIP [44] similarity, and DINOv2 [42] similarity to assess the texture quality of generated assets. Regarding efficiency, we report the inference time cost for synthesizing a single 3D asset on a single A100 GPU. More details will be included in Sec. C.2 of the Appendix.
Baselines. We compare SceneGen with representative 3D scene generation methods, including PartCrafter [34], DepR [72], Gen3DSR [10], and MIDI [23], using their pre-trained models. Specifically, we adopt object masks to specify generation targets for all baselines except for PartCrafter, which does not support mask-based control. Instead, we directly provide PartCrafter with extracted objects and the number of assets as input. Moreover, as PartCrafter and DepR do not offer code for texture rendering, our evaluation of these methods focuses on geometric quality, while visual quality is compared with Gen3DSR and MIDI (relying on MV-Adapter [24] for texture synthesis).
@@ 4.2 Quantitative Results
As presented in Table 1, we draw the following key observations: (i) geometric quality: SceneGen consistently outperforms existing methods across all scene-level and asset-level metrics. This stems from its joint integration of local asset features and global scene context during generation. The interactions among multiple assets facilitate the model in producing physically plausible geometric structures, while the position head further improves the structural realism by explicitly predicting spatial arrangements. (ii) visual quality: SceneGen can render high-quality textures for generated 3D assets without relying on any external texture generation models. Moreover, whether using masked scene images or ground-truth renderings as references, our method consistently achieves the best performance across all metrics. This indicates that our synthesized assets are spatially closer to the ground truth while maintaining superior texture fidelity. and (iii) efficiency: While PartCrafter demonstrates a clear advantage in inference speed, it suffers from limited generation quality and controllability. In contrast, SceneGen achieves both superior quality and a strong balance between quality and efficiency, synthesizing a 3D scene containing four assets with geometry and textures within 2 minutes on a single A100 GPU.
In addition, while the baseline methods, e.g., PartCrafter, DepR, and MIDI have been trained on 3D-FRONT [13], which may overlap with our test data, our SceneGen still consistently outperforms them across all metrics, further demonstrating its effectiveness and superiority.
@@ 4.3 Qualitative Results
Comparisons with baselines. As illustrated in Figure 4, we qualitatively compare SceneGen with existing baselines on both the 3D FUTURE [14] test set and in-the-wild ScnaNet++ [63], where they still struggle with 3D scene generation: PartCrafter lacks controllability over the generated targets and often mistakenly merges distinct assets, while both PartCrafter and DepR are limited to geometry generation and cannot render textures. More critically, all these methods exhibit difficulties in accurately understanding the spatial relationships among assets. In contrast, our proposed SceneGen precisely predicts the spatial relationships among assets and synthesizes multiple 3D assets with accurate geometry and high-quality textures, without relying on any additional tools or optimizations.
Extension to multi-image inputs. Benefiting from our architecture design, SceneGen can seamlessly handle multi-image inputs after being trained exclusively on single-image samples. Given the lack of suitable datasets for quantitative evaluation, we qualitatively assess the impact of multi-image inputs by randomly sampling several scenes from ScanNet++ [63] and employing SAM2 [46] to obtain segmentation masks of corresponding objects. As depicted in Figure 5, compared to single-image inputs, incorporating multi-view images leads to 3D assets with more complete geometry and finer texture details. This illustrates that SceneGen can adaptively integrate complementary information from multiple views to produce higher-quality 3D scenes, further validating its practicality and scalability. More qualitative results will be included in Sec C.1 of the Appendix.
@@ 4.4 Ablation Studies
To validate the efficacy of our modules, we conduct comprehensive evaluations on several variants of SceneGen, assessing both the geometric and visual quality of synthesized scenes. Concretely, we investigate the impact of gradually removing global geometric features ($\mathcal{F}_{\mathrm{global}}^{\mathrm{geo}}$), global visual features ($\mathcal{F}_{\mathrm{global}}^{V}$), mask visual features ($\mathcal{F}_{i}^{\mathrm{mask}}$), as well as substituting the scene-level self-attention block ($\mathcal{A}_{\mathrm{SS}}$) with a simple asset-level self-attention block ($\mathcal{A}_{\mathrm{AS}}$). As depicted in Table 2, we have the following observations: (i) Removing any of the aforementioned components degrades the overall performance, confirming their necessity in SceneGen; (ii) The geometric features primarily affect the structure of synthesized scenes, while the visual features further impact the visual quality; and (iii) The absence of scene-level self-attention blocks eliminates inter-asset interactions during generation, leading to notable performance declines across all metrics. These results strongly demonstrate the necessity and effectiveness of our proposed feature extraction and aggregation modules for SceneGen.
@@ 5 Conclusion
In this paper, we present SceneGen, a novel framework that takes a single scene image and target object masks as input to simultaneously synthesize multiple 3D assets with structure, texture, and relative spatial positions in a single feedforward pass. Specifically, we incorporate dedicated visual and geometric encoders to extract both asset-level and scene-level features, which are effectively fused with our proposed feature aggregation module. Notably, through our meticulous design, SceneGen can naturally generalize to multi-image inputs and achieve even better generation fidelity. Quantitative and qualitative evaluations demonstrate that SceneGen produces physically plausible and mutually consistent 3D assets, significantly outperforming previous methods in terms of generation quality and efficiency.
@@ Appendix A Preliminaries on 3D Foundation Models
TRELLIS. For a 3D asset ($\mathcal{O}$), TRELLIS encodes its geometry and appearance into a unified representation ($\boldsymbol{z}$), denoted as: $\boldsymbol{z}=\{(\boldsymbol{z}_{i},\boldsymbol{p}_{i})\}_{i=1}^{L}$. Here, $\boldsymbol{p}_{i}\in\{0,1,\ldots,D-1\}^{3}$ denotes the positional index of an active voxel in the 3D grid intersecting the surface of $\mathcal{O}$, and $\boldsymbol{z}_{i}\in\mathbb{R}^{C}$ represents the local latent feature attached to the corresponding voxel, with $D$ and $L$ representing the 3D grid resolution and the total number of active voxels, respectively.
@@
VGGT. Trained on large-scale 3D annotated data, VGGT can extract 3D scene features through a purely feedforward network without explicit 3D inductive biases. For single or multi-view RGB inputs ($\{\boldsymbol{I}_{i}\}_{i=1}^{s}$), its aggregator derives scene geometric features ($\{\mathcal{F}_{i}^{\mathrm{geo}}\}_{i=1}^{s}$), represented as:
@@
Here, $\mathcal{F}_{G}^{\mathrm{geo}}$ and $\mathcal{F}_{I}^{\mathrm{geo}}$ denote features extracted by global self-attention and local self-attention layers, respectively. These features are efficiently decoded by lightweight DPT layers [45] into depth maps, point maps, and tracks, validating their rich scene geometric representation capacity.
@@ Appendix B More Details about Training Data
We train SceneGen on the 3D-FUTURE [14] dataset, leveraging its rich textures and diverse lighting conditions to simulate real-world environments and thereby enhance the model’s generalization ability. Additionally, to ensure the model robustly learns the relative spatial relationships among multiple assets, we further scale up training data through data augmentation. Specifically, for a scene with $N$ objects, we iteratively select each asset as the query asset and randomly shuffle the remaining ones during training. Considering GPU memory constraints, we set the maximum number of assets per scene to $N^{\prime}=7$ on a single A100 GPU. For samples containing more than $N^{\prime}$ assets, we randomly select a subset of $N^{\prime}$ assets for training. Furthermore, following TRELLIS [59], we apply its aesthetic score filtering criterion to exclude assets with aesthetic scores below 4.5, thereby ensuring high data quality. The distribution of asset counts across training scenes is illustrated in Figure 6.
@@ Appendix C More Implementation Details
This section provides a holistic explanation of implementation details discussed in the paper. Concretely, Sec. C.1 describes the specific strategies applied to extend SceneGen to multi-image inputs; and Sec. C.2 elaborates on our evaluation protocols.
@@ C.1 Extension to Multi-image Inputs
While SceneGen is primarily designed for 3D scene generation based on a single scene image and trained exclusively on single-view images, it can be seamlessly adapted to multi-view inputs during inference with no need for additional training or fine-tuning. Concretely, during inference, our model can take images of the same scene from multiple viewpoints, along with their corresponding objects and instance masks, as input. Within our SceneGen, the geometric encoder ($\Phi_{G}$) (an off-the-shelf VGGT [53] aggregator) integrates geometric information across different viewpoints to produce better geometric representations for each perspective, thereby enabling SceneGen to synthesize more accurate geometric structures. Finally, we predict the relative positions among different assets from each viewpoint and use the mean of these predictions across all views as the final spatial position output. It is important to note that, to ensure correctness throughout the inference process, the input order of assets and their segmentation masks must remain consistent across all viewpoints.
@@
Table 3: Geometric Metrics Comparisons on Different Point Cloud Alignment Methods.
@@ C.2 Evaluation Protocols
Geometric metrics. Following previous work [23], we conduct geometry evaluation in normalized 3D space (also referred to as canonical space, i.e., $x,y,z\in[-1,1]$), where the ground truth and the synthesized query asset are first rigidly aligned using point cloud registration algorithms. Unlike MIDI [23], which relies on the traditional Iterative Closest Point (ICP [2]) method prone to suboptimal alignment results, we employ FilterReg [17], a faster and more robust point cloud alignment approach. As presented in Table 3, both MIDI and SceneGen achieve better overall performance when aligned via FilterReg, demonstrating the reliability of this alignment method compared to traditional ICP. Moreover, under both alignment strategies, SceneGen consistently outperforms MIDI, indicating that explicitly predicting the spatial positions among assets enables SceneGen to more accurately model the relationships among distinct 3D assets within the scene.
@@ Appendix D More Visualizations
This section presents additional qualitative results on the 3D-FUTURE [14] test set, offering a detailed comparison between our SceneGen and representative baselines. As depicted in Figure 8, we have the following observations: (i) PartCrafter [34] frequently suffers from missing or mixed-up assets due to its inability to control generation via object masks, despite already taking segmented objects and asset counts as input; (ii) Both PartCrafter and DepR [72] can only generate scene geometry without rendering texture details; and (iii) All baseline methods (PartCrafter [34], DepR [72], Gen3DSR [10], and MIDI [23]) share the common limitation of incorrect spatial relationships among synthesized assets. In contrast, our SceneGen fully integrates visual and geometric features within the scene to enable mutual influence among multiple assets during generation, producing 3D scenes with physically plausible geometry and high-quality texture details.
@@ E.1 Limitations
While our SceneGen demonstrates superior performance in 3D scene generation, it is not without its limitations.
Limited to Indoor Generation. While SceneGen demonstrates better texture generation and generalization capabilities compared to previous methods that rely on canonical representations, the narrow training data distribution limits its ability to generalize to non-indoor scenes, restricting its generalization to a broader range of environments.
Asset Collisions and Overlaps. Although SceneGen can generate multiple 3D assets and relative spatial positions in a single feedforward pass, without relying on complex post-processing, it does not always handle contact relationships among objects, occasionally leading to asset overlaps or geometric inconsistencies. This is mainly because our single-stage framework does not explicitly enforce strict spatial or physical constraints among objects.
Reliance on Segmentation Masks. SceneGen inherently requires segmentation masks of the target objects as input. In our current framework, we leverage either ground truth masks or masks pre-extracted by an off-the-shelf SAM 2 [46]. This reliance limits the flexibility of applying SceneGen directly to in-the-wild data to some extent and may potentially result in a lack of robustness against low-quality segmentation masks.